Procedures for constructing a confidence interval for a sample mean are given in section 7-2 on page 319. Example 2 worked on pages 320 and 321 can guide us.

In my homework problem 40, the scenario is data on the salaries of 61 players on a football team. We are given that we are interested in the 95% confidence level and the population standard deviation is 3723 thousand dollars. From section 7-2, we know that to be able to use either the normal or the t​ distribution, either the sample must come from a normally distributed population or the sample size must be greater than 30. In this case, the histogram provided does not seem to have a normal distribution, but the sample size is well above 30.

When the population standard​ deviation is​ known, we use the normal distribution, and thus use zα/2. For 95% confidence level, α = 0.05 and

zα/2 = z0.05/2 = z0.025 = 1.96

Please explain how you would then proceed to determine the 95% confidence interval.
Answer #1

Using normal

To use it we need to have a normal sample and have known population standard deviation. Then using the following formula we can get the confidence interval

(1 - \alpha) confidence interval for population mean

27/0z+1)

Where \bar{x} = Sample Mean

Z_{\alpha/2} = Critical value (found using normal percentage tables or excel function 'normsinv'

\sigma = Population standard deviation

n = Sample size

Margin of error = Z_{\alpha/2} * \frac{\sigma}{\sqrt{n}}

Using t-dist

We use t-dist to approximate the confidence level when we have no population standard deviation and the sample size > 30 (large).

(1 - \alpha) confidence interval for population mean

(ī Fta/2,1-17

Where \bar{x} = Sample Mean

ta/2,1-1 = Critical value (found using t-dist tables or excel function 'tinv')

  S_{x}= Sample standard deviation

  n = Sample size

Margin of error = ta/2,1-1 * \frac{S_{x}}{\sqrt{n}}

We use confidence interval to find a range for the population parameter with (1 - \alpha) certainty that the range will contain the parameter.